OpenStudy (anonymous):

Value Rent-A-Car rents a luxury car at a daily rate of $36.79 plus 10 cents per mile. A business person allotted $90 for car rental each day. How many miles can the business person travel on the $90 ? HELPP PLEASE

OpenStudy (imstuck):

you need to set up an equation given the info you have. The man has 90 to start with, so everything going on with the rental will be equal to that number. The rate o 36.79 is a non-negotiable number that is what the company charges right off the bat. What changes what the man spends is the number of miles he drives, which is what we don't know. WE only know that the miles cost .10 per, right?

OpenStudy (imstuck):

So starting with the $90, we have this so far: 90 = Now we need to figure in the 36.79 that will be charged no matter how many miles are driven that day. Now we have this: 90 = 36.79 + Now we have to figure in the number of miles, x, at 10 cents per mile, which is written as .1x. That's ten cents times an unknown number of miles. So our equation is this: 90 = 36.79 + .1x

OpenStudy (anonymous):

Yes

OpenStudy (imstuck):

You need to solve that for x now to find the number of miles. Can you do that?

OpenStudy (anonymous):

im confused

OpenStudy (cj49):

90-36.79=?

OpenStudy (anonymous):

35.89

OpenStudy (cj49):

x=35.89/0.10=?

OpenStudy (anonymous):

confused again

OpenStudy (cj49):

divide 35.89 by 010

OpenStudy (anonymous):

i get the same answer

OpenStudy (cj49):

wats ur ans

OpenStudy (anonymous):

35.89

OpenStudy (cj49):

35.89/0.10=358.9

OpenStudy (anonymous):

ooo okay. then what do i do ?

OpenStudy (cj49):

so x=358.9

OpenStudy (cj49):

thats ur ans

OpenStudy (anonymous):

it says its wrong

OpenStudy (cj49):

pls recalculate 90-36.79